# DarkAI: Reconstructing the density, velocity and tidal field of dark matter from DESI-like bright galaxy sample

### v1

@@ 4.1 Reconstructing density field
where $P_{\ell}^{\text{rec}}(k)$ is the reconstructed dark-matter power spectrum, $P_{\ell}^{\text{true}}(k)$ is the true dark-matter power spectrum, and $P_{\ell}^{\text{gal}}(k)$ is the redshift-space galaxy density power spectrum. The ratios are computed for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$), providing a measure of the RSD correction performance at different scales.
Figure 10 shows the ratios $R_{\ell}(k)$ for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$) power spectrum. At large scales, the ratios are close to zero, indicating that the RSD correction effectively restores the reconstructed power spectra to the true values. At small scales, the ratios are greater than zero, but typically remain less than 0.5. This suggests that most of the RSDs have been successfully removed, although some residual distortions persist at smaller scales, probably due to less effective reconstruction in high-density regions.
Based on the test results, we conclude that our network, despite using a small sample size, is capable of directly predicting the density field from the galaxy density field and effectively correcting for redshift-space distortions (RSDs). This leads to a reliable reconstruction of the clustering dark-matter density field. However, it is important to note that additional training data is still needed to enhance performance, particularly in high-density regions and on smaller scales.
Figure 12: (Upper panels) Grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slope of the best-fitting relation and the scatter are also indicated in each panel. (Lower panels) Probability distributions of velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$.
Figure 13: Comparison of the velocity-field power spectrum. From left to right, the panels correspond to the velocity magnitude, divergence, and the vorticity, respectively. In each panel, the black solid and blue dashed lines represent the power spectrum of the true and reconstructed velocity field, respectively. The green dot-dashed line shows the cross-correlation power spectra for the velocity magnitude between the true and reconstructed fields.
@@ 4.2 Reconstructing the velocity field
Figure 11 illustrates the comparison between the true and reconstructed dark-matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field, $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field, $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented. The reconstructed field $\vec{v}_{\mathrm{rec}}$ closely matches $\vec{v}_{\mathrm{true}}$, successfully capturing both large-scale coherent flows and small-scale turbulent features. This demonstrates the UNet’s capability to recover the true velocity structure across a wide range of scales. However, the residual map on the right panel highlights discrepancies mainly near the survey boundaries. These deviations are likely caused by boundary effects, as the reconstruction model struggles to fully account for the lack of surrounding information at the edges of the domain. Despite these boundary-related residuals, the overall performance of the reconstruction remains robust, with the primary flow patterns and magnitudes well-reproduced.
@@
The upper panels of Figure 12 show the grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slopes of the best-fit relations are close to unity for both regions, with values around 0.950.950.950.951.001.001.001.00, indicating strong agreement between the reconstructed $v_{\mathrm{rec}}$ and true velocities $v_{\mathrm{true}}$. The scatter values, reflecting the variances from the best-fit lines, are smaller in the center region, demonstrating improved reconstruction precision due to reduced boundary effects. The lower panels of Figure 12 illustrate the probability distributions of velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$. These distributions peak nicely at zero, and the center region shows narrower peaks, further verifying its superior accuracy. In contrast, the broader distributions in the entire survey region underscore the challenges posed by boundary effects and edge distortions. Overall, the results demonstrate the reliability of the reconstruction, with notable improvements in the center survey region compared to the full survey.
@@
To further evaluate the reconstructed velocity field, we calculate the power spectrum, as shown in Figure 13. In each panel, the black solid line represents the power spectrum of the true velocity field, while the blue dashed line corresponds to the reconstructed field. Additionally, the green dot-dashed line shows the cross-correlation power spectra for the velocity magnitude between the true and reconstructed fields, which serves as a measure of their concordance. Overall, the reconstructed velocity field demonstrates strong agreement with the true field across the three components, with some deviations observed.
@@
The cross-power spectrum for the velocity magnitude is systematically lower than unity by approximately 10% on all scales. This discrepancy can be attributed to two factors: (1) survey boundary effects, which truncate the velocity field at the edges of the survey volume and introduce artifacts, particularly on large scales, weakening the correlation between the true and reconstructed modes; and (2) training limitations, as the reconstruction model, typically trained to minimize global error metrics, prioritizes large-scale accuracy at the cost of small-scale fidelity, resulting in suppressed power on smaller scales. For the vorticity power spectrum, the discrepancy at scales $k<0.2\,h\,\mathrm{Mpc}^{-1}$ may arise from the intrinsical nature of the vorticity signal. As vorticity is primarily sourced from small-scale, nonlinear velocity interactions, the reconstruction method may face challenges in accurately reproducing these subtle features at relatively large scale where the signal-to-noise ratio is inherently low. However, at smaller scales $k>0.2\,h\,\mathrm{Mpc}^{-1}$, the reconstructed vorticity shows better agreement with the true field, reflecting the model’s ability to capture localized turbulence. These results underscore the robust overall performance of the reconstruction method, while also highlighting the limitations introduced by survey boundaries and algorithmic challenges.
@@
Figure 14: Eigenvalues of the reconstructed tidal tensor versus those obtained from the true dark-matter density field. The contours for each result encompass 95% and 99% of the grid cells.
@@
Figure 15: Comparison of the classification of the large-scale structure between the true dark-matter density field (left panel) and UNet-reconstructed density field (right panel) in a slice of $1564\times 1564\times 3.05\>h^{-1}{\rm{Mpc}}$. The yellow, yellow-green, green, and black regions are grid cells located at structures classified as clusters, filaments, sheets, and voids, respectively.
@@ 4.3 Reconstructing the tidal field
where $i$ and $j$ are indices taking values 1, 2, or 3, and $\phi$ is the peculiar gravitational potential. We computed $\phi$ from the mass density field using the Poisson equation,
@@
Figure 14 compares the eigenvalues of the tidal field tensor ($T_{1}$, $T_{2}$, $T_{3}$) derived from the UNet-reconstructed dark-matter density field to those obtained from the true dark-matter density field of the mock DESI survey. The relation plots for each eigenvalue reveal a strong alignment along the diagonal, demonstrating excellent agreement between the reconstructed and true tidal fields. The contours, representing 95% and 99% of the grid cells, further confirm that the reconstructed eigenvalues closely match the true values, validating the accuracy of the reconstruction method. This result is consistent with our previous work (Wang et al., 2024), which showed that the UNet-based reconstruction method significantly outperforms traditional halo-based approaches (Wang et al., 2009) in recovering unbiased tidal fields.
The eigenvalues of the tidal field can be used to classify the large-scale structure (LSS) into four categories based on their local morphology:
@@
  • Sheet: grid cells with two negative and one positive eigenvalue;
@@
Figure 15 shows the LSS classifications derived from the true dark-matter density field (left panel) and the UNet-reconstructed density field (right panel). The visual comparison highlights a remarkable consistency between the two panels, with the spatial distributions of structures, including clusters, filaments, sheets and voids, closely matching across the slice. This indicates that the UNet-reconstructed dark-matter density field effectively preserves the topological connectivity and morphological details of the cosmic web. However, small discrepancies are observed, including a slight underestimation of the void volume near the boundaries and misrepresentation of filamentary or sheet regions. These differences likely arise from boundary effects in the survey region, which can truncate the density field near the edges, as well as the inherent challenges in predicting intermediate-density environments, where the contrast between sheets and filaments is less distinct. Despite these minor biases, the UNet reconstruction provides a robust and reliable representation of the tidal field, enabling accurate LSS classification for cosmological studies.
@@ 5 Summary and Conclusion
In this study, we reconstructed the dark matter density, velocity, and tidal fields using a DESI-like Bright Galaxy Sample (BGS) based on a UNet convolutional neural network. Our approach transforms galaxy density fields into the underlying dark matter properties by exploiting the U-Net’s capability to establish precise field-to-field mappings. This method significantly enhances the recovery of large-scale structures and captures the intricate morphology of cosmic web features.
We first generated a DESI-like survey using the Jiutian N-body simulation, which provides high resolution in accordance with Planck2018 cosmological parameters. From the survey, we selected BGS galaxies and the underlying dark matter particles within a redshift range of $0.40.1<z<$ 0.4. The resulting lightcone was carefully trimmed to DESI’s sky coverage of approximately $9625.23\,\mathrm{deg^{2}}$, providing a realistic and comprehensive dataset for both galaxy and dark matter distributions to train and test our models.
We successfully reconstructed the dark matter density field, achieving close agreement with the true field across large-scale structures such as clusters, filaments, and voids. The cross-correlation power spectrum demonstrates a reconstruction accuracy exceeding 98.5% for scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. Moderate discrepancies, with reductions of 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$, were observed in high-density regions due to challenges in reconstructing small-scale features. Furthermore, we evaluated the performance of the RSD correction using multipole moment ratios. These results confirm that the reconstructed field effectively restores the clustering signal by removing most of the RSD effects, achieving strong consistency with the true field, particularly at large scales.
@@
We reconstructed the tidal field by computing the eigenvalues of the tidal tensor and found strong agreement between the reconstructed and true fields. Our approach accurately preserved large-scale structure classifications, including clusters, filaments, sheets, and voids, while maintaining the morphological connectivity of the cosmic web. Minor biases were observed, such as a slight underestimate of the void volume near the boundaries and a misrepresentation of filamentary or sheet regions. These discrepancies are mainly attributed to boundary effects and the challenges of predicting intermediate-density environments. In general, the results confirm the ability of the method to reliably recover the structure of the tidal field.
We have demonstrated that deep learning, specifically a UNet-based architecture, is an effective tool for reconstructing cosmological fields from galaxy samples. Our method consistently achieved high fidelity across density, velocity, and tidal fields, providing a reliable framework for large-scale structure analysis. This study highlights potential applications in observational surveys and enhances our understanding of dark matter distributions and dynamics.

### v2

@@ 4.1 Reconstructing the density field
where $P_{\ell}^{\text{rec}}(k)$ is the reconstructed dark matter power spectrum, $P_{\ell}^{\text{true}}(k)$ is the true dark matter power spectrum, and $P_{\ell}^{\text{gal}}(k)$ is the redshift-space galaxy density power spectra. The ratios are computed for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$), providing a measure of the RSD correction performance at different scales.
Figure 11 shows the ratios $R_{\ell}(k)$ for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$) power spectra. At large scales, the ratios are close to zero, indicating that the RSD correction effectively restores the reconstructed power spectra to the true values. At small scales, the ratios are greater than zero but typically remain less than 0.5. This suggests that most of the RSDs have been successfully removed, although some residual distortions persist at smaller scales, probably due to less effective reconstruction in high-density regions.
Based on the test results, we conclude that our network, despite using a small sample size, is capable of directly predicting the density field from the galaxy density field and effectively correcting for RSDs. This leads to a reliable reconstruction of the clustering mass density field. However, it is important to note that additional training data are still needed to enhance performance, particularly in high-density regions and on smaller scales.
Figure 13: Upper panels: grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in the different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slope of the best-fitting relation and the scatter are also indicated in each panel. Lower panels: probability distributions of the velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$.
Figure 14: Comparison of the velocity field power spectrum. From left to right, the panels correspond to the velocity magnitude, divergence, and the vorticity, respectively. In each panel, the black solid and blue dashed lines represent the power spectra of the true and reconstructed velocity fields, respectively. The green dotted-dashed line shows the cross-correlation power spectra for the velocity magnitude between the true and reconstructed fields.
@@ 4.2 Reconstructing the velocity field
Figure 12 illustrates a comparison between the true and reconstructed dark matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field, $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field, $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented. The reconstructed field $\vec{v}_{\mathrm{rec}}$ closely matches $\vec{v}_{\mathrm{true}}$, successfully capturing both large-scale coherent flows and small-scale turbulent features. This demonstrates UNet’s capability to recover the true velocity structure across a wide range of scales. However, the residual map of the right panel highlights discrepancies mainly near the survey boundaries. These deviations are likely caused by boundary effects, as the reconstruction model struggles to fully account for the lack of surrounding information at the edges of the domain. Despite these boundary-related residuals, the overall performance of the reconstruction remains robust, with the primary flow patterns and magnitudes being well reproduced.
@@
The upper panels of Figure 13 show the grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slopes of the best-fit relations are close to unity for both regions, with values around 0.950.951.001.00, indicating strong agreement between the reconstructed $v_{\mathrm{rec}}$ and true velocities $v_{\mathrm{true}}$. The scatter values, reflecting the variances from the best-fit lines, are smaller in the center region, demonstrating improved reconstruction precision due to reduced boundary effects. The lower panels of Figure 13 illustrate the probability distributions of the velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$. These distributions peak nicely at zero, and the center region shows narrower peaks, further verifying its superior accuracy. In contrast, the broader distributions over the entire survey region underscore the challenges posed by boundary effects and edge distortions. Overall, the results demonstrate the reliability of the reconstruction, with notable improvements in the center survey region compared to the full survey.
@@
To further evaluate the reconstructed velocity field, we calculate the power spectrum, as shown in Figure 14. In each panel, the black solid line represents the power spectrum of the true velocity field, while the blue dashed line corresponds to the reconstructed field. Additionally, the green dotted-dashed line shows the cross-correlation power spectrum for the velocity magnitude between the true and reconstructed fields, which serves as a measure of their concordance. Overall, the reconstructed velocity field demonstrates strong agreement with the true field across the three components, with some deviations observed.
@@
The cross-power spectrum for the velocity magnitude is systematically lower than unity by approximately 10% on all scales. This discrepancy can be attributed to two factors: (1) survey boundary effects, which truncate the velocity field at the edges of the survey volume and introduce artifacts, particularly on large scales, weakening the correlation between the true and reconstructed modes; and (2) training limitations, as the reconstruction model, typically trained to minimize global error metrics, prioritizes large-scale accuracy at the cost of small-scale fidelity, resulting in suppressed power on smaller scales. For the vorticity power spectrum, the discrepancy at scales $k<0.2\,h\,\mathrm{Mpc}^{-1}$ may arise from the intrinsic nature of the vorticity signal. As vorticity is primarily sourced from small-scale, nonlinear velocity interactions, the reconstruction method may face challenges in accurately reproducing these subtle features at relatively large scale where the signal-to-noise ratio is inherently low. However, at smaller scales $k>0.2\,h\,\mathrm{Mpc}^{-1}$, the reconstructed vorticity shows better agreement with the true field, reflecting the model’s ability to capture localized turbulence. These results underscore the robust overall performance of the reconstruction method, while also highlighting the limitations introduced by survey boundaries and algorithmic challenges.
@@
Figure 15: Eigenvalues of the reconstructed tidal tensor versus those obtained from the true mass density field. The contours for each result encompass 95% and 99% of the grid cells.
@@
Figure 16: Comparison of the classification of the LSS between the true mass density field (left panel) and the UNet-reconstructed density field (right panel) in a slice of $1564\times 1564\times 3.05\>h^{-1}{\rm{Mpc}}$. The yellow, yellow-green, green, and black regions show grid cells located at structures classified as clusters, filaments, sheets, and voids, respectively.
Figure 17: Normalized confusion matrix. Each row corresponds to a predicted class derived from the UNet-reconstructed dark matter density field, while each column corresponds to the true class based on the original simulated density field. The matrix entries represent classification probabilities, obtained by normalizing the values within each column such that the total for each true class sums to 1.
@@ 4.3 Reconstructing the tidal field
where $i$ and $j$ are indices taking values 1, 2, or 3, and $\phi$ is the peculiar gravitational potential. We compute $\phi$ from the mass density field using the Poisson equation
@@
Figure 15 compares the eigenvalues of the tidal field tensor ($T_{1}$, $T_{2}$, $T_{3}$) derived from the UNet-reconstructed mass density field to those obtained from the true density field of the mock DESI survey. The relation plots for each eigenvalue reveal a strong alignment along the diagonal, demonstrating excellent agreement between the reconstructed and true tidal fields. The contours, representing 95% and 99% of the grid cells, further confirm that the reconstructed eigenvalues closely match the true values, validating the accuracy of the reconstruction method. This result is consistent with our previous work (Wang24), which showed that the UNet-based reconstruction method significantly outperforms traditional halo-based approaches (Wang et al., 2009) in recovering unbiased tidal fields.
The eigenvalues of the tidal field can be used to classify the LSS into four categories based on their local morphology:
@@
  • Sheet: grid cells with two negative and one positive eigenvalues;
@@
Figure 16 shows the LSS classifications derived from the true mass density field (left panel) and the UNet-reconstructed density field (right panel). The visual comparison highlights a remarkable consistency between the two panels, with the spatial distributions of structures, including clusters, filaments, sheets and voids, closely matching across the slice. This indicates that the UNet-reconstructed density field effectively preserves the topological connectivity and morphological details of the cosmic web.
Figure 17 presents the normalized confusion matrix evaluating the classification performance of the UNet-reconstructed LSS field. The matrix compares the predicted LSS types derived from the reconstructed dark matter density field against the ground-truth classification from the original simulation. Each row corresponds to the predicted class, while each column corresponds to the true class, and the matrix entries represent the column-normalized classification probabilities. The diagonal elements indicate the proportion of correctly predicted classes, highlighting that sheets and filaments are reconstructed with relatively high fidelity (79.56% and 77.64%, respectively), followed by voids (72.09%) and clusters (63.26%). Clusters are frequently misclassified as filaments (35.37%), reflecting the network’s tendency to underestimate high-density peaks due to their sparsity and sharp gradients. Voids, while generally well recovered, show moderate confusion with sheets (27.02%), likely due to boundary smoothing in the reconstructed field, which reduces the contrast between underdense void interiors and adjacent sheet-like structures. Overall, the confusion matrix confirms that the UNet model robustly reconstructs the cosmic web morphology, especially in filamentary and sheet-like environments, while revealing relatively higher level of challenge in terms of reconstructing dense cluster regions.
5 Discussion
The reconstruction framework presented in this work demonstrates robust performance in recovering the overall dark matter density and velocity fields. However, it exhibits scale- and environment-dependent limitations that that may affect its applicability to certain scientific analyses.
One such limitation arises in the reconstruction of high-density peaks, where the model systematically underestimates the true density in regions with $1+\delta>7$. Although such regions constitute only about 0.2% of the total grid volume, they correspond to massive cluster-scale structures that host significant cosmological information. This underestimation likely stems from the rarity of these regions in the training data and the intrinsic limitations of convolutional networks in capturing sharp density contrasts. As a result, analyses that require accurate reconstructions of cluster morphology or mass, such as environmental dependence studies (see Section 4.3), non-Gaussianity analyses (e.g., Shandera et al., 2013), gravitational lensing (e.g., Hoekstra et al., 2013), and Sunyaev–Zel’dovich cluster catalogs (e.g., Hilton et al., 2021), can be adversely affected.
In addition, the power spectrum analysis reveals a scale-dependent discrepancy at $k>0.1~h~\mathrm{Mpc}^{-1}$, where the reconstructed field underestimates the true power by up to ahout 20%. This deviation, consistent with the suppressed densities in high-density regions, indicates a limitation in recovering nonlinear clustering information. While the reconstruction remains reliable for LSS statistics, it biases the small-scale power spectrum, limiting its utility in testing nonlinear gravitational models (e.g., Cataneo et al., 2019) and quantifying baryonic feedback effects(e.g., Preston et al., 2024). Inaccuracies in the power spectrum can also compromise the accuracy of cosmological parameter inference based on small-scale power(e.g., Oddo et al., 2021). Additional consequences may include reduced fidelity in reconstructing initial conditions or in generating forward-modeled observables (e.g., Wang et al., 2016).
The reconstructed velocity field also exhibits a systematic suppression in amplitude across all scales, with a roughly 10% deficit in the cross-power spectrum relative to the true velocity field. This underestimation can impact a range of scientific applications that rely on accurate 3D velocity information. For example, kinetic Sunyaev–Zel’dovich tomography, which cross-correlates reconstructed velocity fields with cosmic microwave background temperature maps(Shao et al., 2011), requires the faithful recovery of both amplitude and direction to avoid biasing the inferred baryon distribution. Similarly, the use of velocity fields to trace bulk flows (Boruah et al., 2020), analyze void dynamics(Ceccarelli et al., 2016), or reconstruct initial conditions(Doumler et al., 2013) depends on preserving coherent structures across a wide range of scales. The suppressed velocity power may therefore reduce the effectiveness of these analyses. Improving velocity reconstruction may require velocity-specific loss functions and enhanced boundary treatment to support precision applications in velocity-based cosmology.
Therefore, the current reconstruction framework performs well in capturing large-scale and intermediate-density structures. However, its accuracy degrades in regimes characterized by nonlinear structures and velocity complexity. Extending its applicability to these domains will likely require improvements in the training methodology and model architecture, including adaptive loss functions, focused data augmentation, and multiscale designs to better capture nonlinear structure and velocity information. We will pursue these enhancements in future work to broaden the framework’s scientific utility.
@@ 6 Summary and Conclusion
In this study, we have reconstructed the MTV fields using a DESI-like BGS based on a UNet convolutional neural network. Our approach transforms galaxy density fields into the underlying dark matter properties by exploiting UNet’s capability to establish precise field-to-field mappings. This method significantly enhances the recovery of LSSs and captures the intricate morphology of cosmic web features.
We first generated a DESI-like survey using the Jiutian N-body simulation, which provides high resolution in accordance with Planck2018 cosmological parameters. From the survey, we selected BGS galaxies and underlying dark matter particles within a redshift range of $0.1<z<0.4$. The resulting light cone was carefully trimmed to DESI’s sky coverage of approximately $9625.23\,\mathrm{deg^{2}}$, providing a realistic and comprehensive dataset for both galaxy and dark matter distributions to train and test our models.
We successfully reconstructed the mass density field, achieving close agreement with the true field across LSSs such as clusters, filaments, and voids. The cross-correlation power spectrum demonstrates a reconstruction accuracy exceeding 98.5% for scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. Moderate discrepancies, with reductions of 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$, were observed in high-density regions due to challenges in reconstructing small-scale features. Furthermore, we evaluated the performance of the RSD correction using multipole moment ratios. These results confirm that the reconstructed field effectively restores the clustering signal by removing most of the RSD effects, achieving strong consistency with the true field, particularly at large scales.
@@
We reconstructed the tidal field by computing the eigenvalues of the tidal tensor and found strong agreement between the reconstructed and true fields. Our approach accurately preserved LSS classifications, including clusters, filaments, sheets, and voids, while maintaining the morphological connectivity of the cosmic web. Minor biases were observed, such as a slight underestimate of the void volume near the boundaries and a misrepresentation of filamentary or sheet regions. These discrepancies are mainly attributed to boundary effects and the challenges of predicting intermediate-density environments. In general, the results confirm the ability of the method to reliably recover the structure of the tidal field.
We have shown that a UNet-based deep learning framework can effectively reconstruct MTV fields from galaxy samples. While the method performs well overall, its accuracy declines in high-density regions and shows moderate suppression in velocity field reconstruction. These limitations point to directions for future improvement. Nonetheless, the approach provides a promising tool for LSS analysis in forthcoming observational surveys.
7 Data AVAILABILITY
The dataset of the 3D fields of the simulated DESI BGS galaxy and dark matter distributions will be made available through the PaperData Repository of the National Astronomical Data Center of China: https://doi.org/10.12149/101654.
8 Acknowledgments